Circular No. 7677

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

WZ SAGITTAE P. J. Wheatley, University of Leicester; E. Kuulkers, Space Research

Organization Netherlands (SRON) and Utrecht University; J. J. Drake, Smithsonian Astrophysical Observatory; J. S. Kaastra, SRON; C. W. Mauche, Lawrence Livermore National Laboratory; S. G. Starrfield, Arizona State University; and R. M. Wagner, University of Arizona; together with a team from the Chandra X-ray Center, report: “The Chandra X-ray Observatory made a 20000-s low-energy-transmission-grating (LETG) observation of WZ Sge beginning on July 27.11 UT, three and a half days into outburst (IAUC 7669, 7670). The mean zeroth-order count rate was 4.5 counts/s, and this is modulated on the orbital period of WZ Sge with a double-peaked profile. The dispersed spectrum extends from 0.2 to 17.5 nm and has a mean count rate of 20 counts/s. Longward of 6.5 nm, the spectrum is dominated by a forest of broad (FWHM between 800 and 1200 km/s) emission lines of such species as O V–VI, Ne V–VIII, Mg V–VII, and Fe VII–IX. Shortward of 6.5 nm, weaker lines of highly ionized ions are seen, with O VIII (1.89 nm) the strongest. The long-wavelength LETG spectrum is very similar to the EUVE spectrum of OY Car in superoutburst (Mauche and Raymond 2001, Ap.J. 541, 924) and appears to be formed by the same mechanism — scattering of boundary-layer radiation in the system’s accretion-disk wind. A second Chandra observation of WZ Sge was made with the ACIS-S CCD detector beginning July 29.70 and lasting 5000 s. The source is detected at x-ray energies up to 7 keV, and the x-ray spectrum is dominated also by strong emission lines. Most features can be accounted for with a multi-temperature thermal plasma model with the exception of an anomalously strong emission line at 2.4 keV, probably S XV (He-like). Five additional Chandra observations of WZ Sge are planned, the first on Aug. 6 and the last in the week beginning Sept. 17.”

V2274 CYGNI J. C. Wilson, Cornell University; and R. J. Rudy and D. K. Lynch, The

Aerospace Corporation, report 0.8–2.5-μm spectrophotometry of V2274 Cyg (= N Cyg 2001) using the Palomar 1.5-m telescope (+ CorMASS) on July 31.3 UT, when the nova was in early decline: “The spectrum showed narrow, low-excitation lines that were typical of the early emission stage. The H I Paschen and Brackett lines were present, along with He I 1.0831μm, which was weak. Also present were the Lyβ-fluoresced O I lines and possibly permitted lines of C I and N I.”

2001 August 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green